# Rethinking Machine Unlearning for Large Language Models

### v4

@@ 1 Introduction
Large language models (LLMs) have shown exceptional proficiency in generating text that closely resembles human-authored content. However, their ability to memorize extensive corpora may also lead to ethical and security concerns. These include societal biases and stereotyping Bender et al. (2021); Motoki et al. (2023); Kotek et al. (2023), the generation of sensitive, private, harmful, or illegal content Nasr et al. (2023); Wen et al. (2023); Karamolegkou et al. (2023); Patil et al. (2024), ease of jailbreaking Wei et al. (2023); Zou et al. (2023); Liu et al. (2023b), and possible malicious use in developing cyberattacks or bioweapons Barrett et al. (2023); Hendrycks et al. (2023); Li et al. (2024a). These concerns emphasize the need to adeptly and efficiently tailor pre-trained LLMs to suit diverse safety contexts while meeting specific requirements of users and sectors.
With the costly and prolonged training periods of LLMs, retraining these models to eliminate undesirable data effects is often impractical Brown et al. (2020); Yao et al. (2024). Machine unlearning (MU) has emerged as an alternative to remove the influence of undesirable data and associated model capabilities from the pre-trained models Cao & Yang (2015); Bourtoule et al. (2021); Nguyen et al. (2022); Si et al. (2023); Zhang et al. (2023a); Eldan & Russinovich (2023); Yao et al. (2023). For example, MU is used as a strategy to prevent the generation of copyrighted material from the Harry Potter series Eldan & Russinovich (2023). In the context of classification tasks, MU has been extensively studied Ginart et al. (2019); Neel et al. (2021); Ullah et al. (2021); Sekhari et al. (2021); Golatkar et al. (2020); Jia et al. (2023). Yet, its application and understanding in LLMs remains limited, where models are typically used for generative tasks such as summarization, sentence completion, paraphrasing, and question answering. Therefore, this paper specifically concentrates on exploring the MU problems in LLMs, termed ‘LLM Unlearning’.
As data-model scales continue to grow, the emergence of LLM unlearning introduces new challenges and complexities, as will be elaborated on in Sec. 2. For example, current research efforts Lu et al. (2022); Jang et al. (2022); Ilharco et al. (2022); Eldan & Russinovich (2023); Wu et al. (2023b); Yu et al. (2023); Zhang et al. (2023c); Yao et al. (2023) suffer from a lack of standardized corpora and principled evaluation for LLM unlearning. Although preliminary surveys of LLM unlearning have been provided in Si et al. (2023); Zhang et al. (2023a), this paper is, to the best of our knowledge, the first to offer a thorough and in-depth review of LLM unlearning. The key contributions are summarized below. Fig. 1 provides an overview of the LLM unlearning landscape that we explore.
@@ 2 Related Work
LLM unlearning has garnered attention for addressing trustworthiness concerns such as toxicity Lu et al. (2022), copyright and privacy Jang et al. (2022); Eldan & Russinovich (2023); Wu et al. (2023b), fairness Yu et al. (2023), hallucination Yao et al. (2023), malicious usage Li et al. (2024a), and sensitive knowledge Barrett et al. (2023); Hendrycks et al. (2023). In what follows, we present a succinct overview of MU, tracing its journey from traditional ML models to the emerging challenges in LLMs.
@@
The study of MU can be traced back to non-LLMs in response to data protection regulations such as ‘the right to be forgotten’ Cao & Yang (2015); Hoofnagle et al. (2019); Bourtoule et al. (2021); Nguyen et al. (2022). Due to its capability of assessing data influence on model performance, the landscape of MU has expanded to encompass diverse domains, such as image classification Ginart et al. (2019); Golatkar et al. (2020); Neel et al. (2021); Ullah et al. (2021); Sekhari et al. (2021), text-to-image generation Gandikota et al. (2023); Zhang et al. (2023b); Kumari et al. (2023); Fan et al. (2024b), federated learning Liu et al. (2020); Wang et al. (2022); Che et al. (2023); Liu et al. (2023c); Halimi et al. (2022), graph neural networks Chen et al. (2022b); Chien et al. (2022); Wu et al. (2023a), and recommendation Sachdeva et al. (2024); Chen et al. (2022a); Xu et al. (2023); Li et al. (2024b); Wang et al. (2024b).
In the literature, ‘exact’ unlearning, which involves retraining the model from scratch after removing specific training data points, is often considered the gold standard. However, this approach comes with significant computational demands and requires access to the entire training set Thudi et al. (2022a). To address these challenges, many research efforts have shifted towards the development of scalable and effective approximate unlearning methods Golatkar et al. (2020); Warnecke et al. (2021); Becker & Liebig (2022); Thudi et al. (2022a); Jia et al. (2023); Chen et al. (2023a). In addition, probabilistic methods with certain provable removal guarantees have been explored, often leveraging the concept of differential privacy Ginart et al. (2019); Guo et al. (2019); Neel et al. (2021); Ullah et al. (2021); Sekhari et al. (2021).
@@
LLM unlearning introduces new challenges and complexities. First, LLMs are trained on massive amounts of data, which can unintentionally introduce biases and the memorization of personal and confidential information. Accordingly, it becomes challenging to precisely define and localize the ‘unlearning targets’, such as the subset of the training set or a knowledge concept that needs to be removed. Therefore, current studies on LLM unlearning Lu et al. (2022); Jang et al. (2022); Ilharco et al. (2022); Eldan & Russinovich (2023); Wu et al. (2023b); Yu et al. (2023); Zhang et al. (2023c); Yao et al. (2023) are typically context and task-dependent. There is a lack of standardized corpora for LLM unlearning. Second, the growing size of LLMs and the rise of black-box access to LLM-as-a-service present challenges in developing scalable and adaptable MU techniques to LLMs Bucknall & Trager (2023); Casper et al. (2024a). This also affects performance evaluation, given the absence of retraining as a benchmark. To address these challenges, previous studies have proposed approaches like in-context unlearning Pawelczyk et al. (2023) and fictitious unlearning Maini et al. (2024), where the former enables unlearning on black-box models, and the latter provides a synthetic for ease of retraining. Third, the scope of unlearning is often underspecified for LLMs. This issue is similar to challenges faced in model editing Mitchell et al. (2022). For instance, effective unlearning should ensure that LLMs delete knowledge of the targeted data within the predefined scope while simultaneously maintaining its utility for data outside of this scope. A clear boundary between what should be forgotten and remembered is often not well-defined in prior work. Fourth, despite the potential of LLM unlearning in diverse applications, there is a notable absence of comprehensive and reliable evaluation. For example, recent studies Shi et al. (2023); Patil et al. (2024); Lynch et al. (2024); Zhang et al. (2024a) have demonstrated that sensitive information can be reverse-engineered from an LLM post-unlearning, even if efforts were made to delete this information. This highlights the need for thorough and adversarial evaluations and the design of more mechanistic methods to guarantee the authenticity of unlearning.

### v5

@@ 1 Introduction
Large language models (LLMs) have shown exceptional proficiency in generating text that closely resembles human-authored content. However, their ability to memorize extensive corpora may also lead to ethical and security concerns. These include societal biases and stereotyping (Bender et al., 2021; Motoki et al., 2023; Kotek et al., 2023), the generation of sensitive, private, harmful, or illegal content (Nasr et al., 2023; Wen et al., 2023; Karamolegkou et al., 2023; Patil et al., 2024), ease of jailbreaking (Wei et al., 2023; Zou et al., 2023; Liu et al., 2023b), and possible malicious use in developing cyberattacks or bioweapons (Barrett et al., 2023; Hendrycks et al., 2023; Li et al., 2024a). These concerns emphasize the need to adeptly and efficiently tailor pre-trained LLMs to suit diverse safety contexts while meeting specific requirements of users and sectors.
With the costly and prolonged training periods of LLMs, retraining these models to eliminate undesirable data effects is often impractical (Brown et al., 2020; Yao et al., 2024). Machine unlearning (MU) has emerged as an alternative to remove the influence of undesirable data and associated model capabilities from the pre-trained models (Cao & Yang, 2015; Bourtoule et al., 2021; Nguyen et al., 2022; Si et al., 2023; Zhang et al., 2023a; Eldan & Russinovich, 2023; Yao et al., 2023). For example, MU is used as a strategy to prevent the generation of copyrighted material from the Harry Potter series (Eldan & Russinovich, 2023). In the context of classification tasks, MU has been extensively studied (Ginart et al., 2019; Neel et al., 2021; Ullah et al., 2021; Sekhari et al., 2021; Golatkar et al., 2020; Jia et al., 2023). Yet, its application and understanding in LLMs remains limited, where models are typically used for generative tasks such as summarization, sentence completion, paraphrasing, and question answering. Therefore, this paper specifically concentrates on exploring the MU problems in LLMs, termed ‘LLM Unlearning’.
As data-model scales continue to grow, the emergence of LLM unlearning introduces new challenges and complexities, as will be elaborated on in Sec. 2. For example, current research efforts (Lu et al., 2022; Jang et al., 2022; Ilharco et al., 2022; Eldan & Russinovich, 2023; Wu et al., 2023b; Yu et al., 2023; Zhang et al., 2023c; Yao et al., 2023) suffer from a lack of standardized corpora and principled evaluation for LLM unlearning. Although preliminary surveys of LLM unlearning have been provided in (Si et al., 2023; Zhang et al., 2023a), this paper is, to the best of our knowledge, the first to offer a thorough and in-depth review of LLM unlearning. The key contributions are summarized below. Fig. 1 provides an overview of the LLM unlearning landscape that we explore.
@@ 2 Related Work
LLM unlearning has garnered attention for addressing trustworthiness concerns such as toxicity (Lu et al., 2022), copyright and privacy (Jang et al., 2022; Eldan & Russinovich, 2023; Wu et al., 2023b), fairness (Yu et al., 2023), hallucination (Yao et al., 2023), malicious usage (Li et al., 2024a), and sensitive knowledge (Barrett et al., 2023; Hendrycks et al., 2023). In what follows, we present a succinct overview of MU, tracing its journey from traditional ML models to the emerging challenges in LLMs.
@@
The study of MU can be traced back to non-LLMs in response to data protection regulations such as ‘the right to be forgotten’ (Cao & Yang, 2015; Hoofnagle et al., 2019; Bourtoule et al., 2021; Nguyen et al., 2022). Due to its capability of assessing data influence on model performance, the landscape of MU has expanded to encompass diverse domains, such as image classification (Ginart et al., 2019; Golatkar et al., 2020; Neel et al., 2021; Ullah et al., 2021; Sekhari et al., 2021), text-to-image generation (Gandikota et al., 2023; Zhang et al., 2023b; Kumari et al., 2023; Fan et al., 2024b), federated learning (Liu et al., 2020; Wang et al., 2022; Che et al., 2023; Liu et al., 2023c; Halimi et al., 2022), graph neural networks (Chen et al., 2022b; Chien et al., 2022; Wu et al., 2023a), and recommendation (Sachdeva et al., 2024; Chen et al., 2022a; Xu et al., 2023; Li et al., 2024b; Wang et al., 2024b).
In the literature, ‘exact’ unlearning, which involves retraining the model from scratch after removing specific training data points, is often considered the gold standard. However, this approach comes with significant computational demands and requires access to the entire training set (Thudi et al., 2022a). To address these challenges, many research efforts have shifted towards the development of scalable and effective approximate unlearning methods (Golatkar et al., 2020; Warnecke et al., 2021; Becker & Liebig, 2022; Thudi et al., 2022a; Jia et al., 2023; Chen et al., 2023a). In addition, probabilistic methods with certain provable removal guarantees have been explored, often leveraging the concept of differential privacy (Ginart et al., 2019; Guo et al., 2019; Neel et al., 2021; Ullah et al., 2021; Sekhari et al., 2021).
@@
LLM unlearning introduces new challenges and complexities. First, LLMs are trained on massive amounts of data, which can unintentionally introduce biases and the memorization of personal and confidential information. Accordingly, it becomes challenging to precisely define and localize the ‘unlearning targets’, such as the subset of the training set or a knowledge concept that needs to be removed. Therefore, current studies on LLM unlearning (Lu et al., 2022; Jang et al., 2022; Ilharco et al., 2022; Eldan & Russinovich, 2023; Wu et al., 2023b; Yu et al., 2023; Zhang et al., 2023c; Yao et al., 2023) are typically context and task-dependent. There is a lack of standardized corpora for LLM unlearning. Second, the growing size of LLMs and the rise of black-box access to LLM-as-a-service present challenges in developing scalable and adaptable MU techniques to LLMs (Bucknall & Trager, 2023; Casper et al., 2024a). This also affects performance evaluation, given the absence of retraining as a benchmark. To address these challenges, previous studies have proposed approaches like in-context unlearning (Pawelczyk et al., 2023) and fictitious unlearning (Maini et al., 2024), where the former enables unlearning on black-box models, and the latter provides a synthetic for ease of retraining. Third, the scope of unlearning is often underspecified for LLMs. This issue is similar to challenges faced in model editing (Mitchell et al., 2022). For instance, effective unlearning should ensure that LLMs delete knowledge of the targeted data within the predefined scope while simultaneously maintaining its utility for data outside of this scope. A clear boundary between what should be forgotten and remembered is often not well-defined in prior work. Fourth, despite the potential of LLM unlearning in diverse applications, there is a notable absence of comprehensive and reliable evaluation. For example, recent studies (Shi et al., 2023; Patil et al., 2024; Lynch et al., 2024; Zhang et al., 2024a) have demonstrated that sensitive information can be reverse-engineered from an LLM post-unlearning, even if efforts were made to delete this information. This highlights the need for thorough and adversarial evaluations and the design of more mechanistic methods to guarantee the authenticity of unlearning.